• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 12:45:36 on September25,2005

    Entry number 155898

    keyword=Compton cavity not locked

    Find that Compton was off, DAQ hangup; did a CODA reboot and
    restart. DAQ comes back up OK, and new run starts OK, but cavity power
    is zero. Laser is on, but cavity is not locked. Try the cavity lock
    instructions: Servo ON, slow Ramp procedurr. No joy. Try a few
    more times, adjusting the "tracking level" down to 0.3 as per
    instructions. Still no joy. Page Sirish, but no answer yet. 
    
    We haven't had Comapton data since about 5:00 AM.
    
    What is next to try?